# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

### v1

@@
Abstract
With the advent of large language models (LLMs), numerous software service providers (SSPs) are dedicated to developing LLMs customized for code generation tasks, such as CodeLlama and Copilot. However, these LLMs can be leveraged by attackers to create malicious software, which may pose potential threats to the software ecosystem. For example, they can automate the creation of advanced phishing malware. To address this issue, we first conduct an empirical study and design a prompt dataset, MCGTest, which involves approximately $400$ person-hours of work and consists of $406$ malicious code generation tasks. Utilizing this dataset, we propose MCGMark, the first robust, code structure-aware, and encodable watermarking approach to trace LLM-generated code. We embed encodable information by controlling the token selection and ensuring the output quality based on probabilistic outliers. Additionally, we enhance the robustness of the watermark by considering the structural features of malicious code, preventing the embedding of the watermark in easily modified positions, such as comments. We validate the effectiveness and robustness of MCGMark on the DeepSeek-Coder. MCGMark achieves an embedding success rate of $88.9\%$ within a maximum output limit of $400$ tokens. Furthermore, it also demonstrates strong robustness and has minimal impact on the quality of the output code. Our approach assists SSPs in tracing and holding responsible parties accountable for malicious code generated by LLMs.
Index Terms:
@@ 1 Introduction
Code generation has become an important topic in software engineering [1], enabling the automatic generation of code snippets from the natural language requirements provided by users. This advancement significantly reduces the manual coding effort and simplifies the software development process [2]. Recently, with the advent and development of large language models (LLMs), their capacity for enhancing code-related tasks has been notably recognized [3]. Meanwhile, Software Service Providers (SSP) are dedicating efforts to develop LLMs specifically tailored for code generation tasks, such as CodeLlama [4] and DeepSeek-Coder [5].
Although LLMs offer remarkable capabilities for code generation, they are often exploited by attackers for misbehaving activities. Previous research [6] reveals that participants in online criminal forums are exploiting LLMs to develop malicious software, e.g., spyware and ransomware. Furthermore, reports from various organizations, such as Meta [7] and Cybersecurity [8], indicate a rising trend in malicious software and network attacks caused by LLMs. Numerous instances and posts also demonstrate the use of LLMs to generate malicious code on various technical forums [9, 10], which could potentially endanger users’ assets.
However, SSPs face significant challenges in tracing the sources of malicious code generated by LLMs [11]. Identifying the specific malicious developer who utilizes an LLM to generate a given piece of malicious code remains an unresolved challenge in the current landscape. This makes the malicious use of LLMs for code generation a low-risk but high-return activity. To mitigate this issue, SSPs like OpenAI have also proposed various solutions, such as zero-shot detectors [12] and fine-tuning language model detectors [13]. However, subsequent research has shown that these solutions are ineffective or have poor robustness [14]. For example, OpenAI’s classifier, which had an accuracy rate of only $26\%$, has now been discontinued by OpenAI [15].
Watermarking technology is a promising solution for tracing the origin of the content generated by LLMs [16, 17]. It involves incorporating explicit or hidden features, such as special lexical substitution, into the content to distinguish whether it is generated by LLMs or humans [18]. However, current watermarking methods may not meet SSP’s requirements for tracing the source of malicious code [19, 20]. Firstly, they suffer from poor traceability and implicitness. The most advanced LLM watermarking technologies can only identify whether a piece of text was generated by an LLM, but they cannot reflect the user’s identity [21, 22]. This limitation prevents SSP from tracing malicious developers. Additionally, most watermarking techniques are post-processing based [23, 24, 17], typically involving word substitution and statistical rules, which only modify the LLM’s generated output. Such watermarks often leave noticeable characteristics in the code, making them more easily detectable and modifiable by malicious developers [16]. Secondly, maintaining code generation quality is challenging. The most advanced implicit watermarking techniques primarily target text generation [25, 26], while code typically has a more rigid structure and stricter syntax rules [27]. Embedding implicit watermarks during code generation while preserving code quality presents a significant challenge. Lastly, robustness is a concern. Certain elements in the code, such as comments and variable names, can be modified without affecting the code’s functionality [28]. Malicious developers can easily alter these elements, making the watermark vulnerable to removal. Therefore, it is necessary to consider the structural characteristics of the code when embedding watermarks to ensure that they are not easily compromised.
@@
To address the aforementioned issues, in this paper, we propose MCGMark, the first watermarking method specialized for tracing code generated by LLMs. Our approach implicitly embeds encodable watermarks during the code generation process, considering the code’s structure and ensuring the quality of the generated code. Firstly, MCGMark implicitly embeds encodable watermarks by controlling the LLM’s token selection, ensuring the watermark is difficult to discern while reflecting the generator’s identity. Secondly, MCGMark maintains the output quality of the LLM by handling the probabilistic outliers of candidate tokens. Lastly, we design a watermark skipping mechanism based on code structure and syntax rules, allowing MCGMark to decide whether to embed the watermark in subsequent code elements based on the code already generated. This ensures that watermarks are not embedded in easily modifiable code elements, such as comments and variable names, thereby enhancing the robustness of the watermark. Additionally, we conduct an empirical study on existing instances of malicious code. We collect $129$ real malicious code examples generated by LLMs and analyze $21,959$ malicious code repositories on GitHub. Based on the empirical study, we construct the first prompt dataset for LLM-generated malicious code, MCGTest, consisting of $406$ malicious code generation tasks to inspire watermark design and test its effectiveness.
MCGMark is designed as a plug-in, eliminating the need for additional models and data during both watermark embedding and detection processes. To implement MCGMark, the SSP only needs to adapt its token-matching rules to align with the specific LLM vocabulary in use. We adapt and apply MCGMark on DeepSeek-Coder-6.7b-instruct to evaluate its performance. MCGMark embed a total 24-bit watermark in $400$ tokens, achieving a watermark embedding success rate of $88.9\%$. Additionally, we conduct a user study involving $500$ program pairs and $1200$ modification attacks. The analysis demonstrates MCGMark’s effectiveness in maintaining code generation quality and resisting modification attacks.
@@
  • We construct the first malicious prompt dataset, MCGTest, consisting of $406$ prompts for LLM’s malicious code generation. based on MCGTest, we design and test the proposed watermarking method. This dataset also inspires further adversarial work against LLM’s malicious code generation.
@@
  • We propose MCGMark, the first robust, code structure-aware, and encodable watermarking scheme to trace LLM-generated code. MCGMark implicitly embeds user identity information in code generation, ensuring both code quality and robustness against watermark tampering.
  • We apply MCGMark to DeepSeek-Coder-6.7b-instruct. The experimental results show that it successfully embeds a $24$-bit watermark and achieves a success rate of $88.9\%$ under a maximum LLM output of $400$ tokens. Additionally, MCGMark ensures code generation quality and withstands various modification attacks.
  • We publish the source code of MCGMark, and MCGTest at: https://github.com/KevinHeiwa/MCGTM
2 Background and Challenges
Figure 1: The motivation example and challenges of design watermark against LLM-generated malicious code.
2.1 Code Generation of LLM
@@
Large Language Models (LLMs) are large-scale language models based on the Transformer architecture [29] and are trained on massive corpora, typically with billions of parameters or more [30]. In recent research, LLMs have demonstrated impressive performance in code generation tasks [31] , which can significantly improve software development efficiency [32]. The performance of LLMs in code generation tasks has received extensive research attention [33, 32]. Moreover, some Software Service Providers (SSPs) have trained/fine-tuned dedicated LLMs specifically for code generation, known as Code LLMs. Currently, SSPs have developed numerous popular Code LLMs, such as Code Llama [4], DeepSeek-Coder [5], and StarCoder2 [34].
2.2 Motivation
@@
LLMs are increasingly being exploited by malicious developers for generating malicious code [35, 36, 37]. Numerous instances have demonstrated the high efficacy of LLMs in producing harmful software. Fig. 1 illustrates a real-world example where an LLM was prompted to generate code for stealing browser history  [^1] https://github.com/AI-Generated-Scripts/GPT-Malware. This irresponsible utilization of LLMs for malicious code generation could pose a significant threat to the security of the software ecosystem.
However, unlike malicious text, which often exhibits overt harmful features, certain code functionalities (e.g., keyboard logging) may not exhibit apparent malicious characteristics [38]. This subtlety renders LLMs vulnerable to exploitation for generating harmful code [39]. Furthermore, malicious developers can employ instruction hijacking [40] and jailbreaking [41] to further facilitate the generation of malicious code through LLMs. Therefore, it is imperative to design alternative approaches for LLMs to combat malicious code generation, with watermarking schemes emerging as one of the most promising solutions [20].
2.3 Challenges
@@
Designing watermarks to trace the generation of malicious code introduces several novel challenges.
  • Traceability & Implicitness. The watermark must be implicit and accurately reflect the user’s ID. Fig.1.(1).a shows an example of a watermark from the literature [21], but applying this watermark only indicates whether the code was generated by an LLM, failing to trace back to a specific user. Fig.1. (1).b illustrates a pattern from the watermarking technique in the literature [23], which is a post-processing watermark. This technique does not participate in code generation, it processes the code after generation using rules such as code transformation [24]. However, this watermark depends on predefined code transformation patterns, limiting its adaptability and failing to ensure compatibility with all code variations. Furthermore, relying on predefined patterns makes the watermarked code exhibit noticeable characteristics, which can be easily detected and disrupted by malicious developers.
@@
  • Ensuring Code Generation Quality. The embedding of watermarks must maintain the quality of the generated code. Fig.1.(2) shows an example of a hard watermark from the literature [25]. In this instance, watermark embedding significantly degraded the code generation quality, rendering the LLM-generated code unusable. Compared to text, code contexts typically exhibit lower entropy and follow more fixed patterns in generation [27]. Some multi-encoding watermarking techniques rely on the strong generative capabilities of LLMs to mitigate the impact of watermarking on generation quality [42, 43]. However, this approach is not suitable for code generation. Even minor perturbations in code can significantly affect its functionality and usability. Thus, during the LLM’s code generation process, minimizing the impact of watermark embedding on code quality is a crucial challenge.
@@
  • Resistant to Tampering. Code elements, such as comments and variable names, can be altered without affecting the code’s functionality. If a watermark is embedded in these elements, it can be easily tampered with. Fig.1.(3).a demonstrates a watermarking scheme from the literature [19], where the watermark is added to variable names and can be easily modified. Fig.1. (3).b shows an example of a soft watermark from the literature [25], where the watermark is added to comments and can also be easily disrupted. However, during the online embedding of watermarks, the generation of code and the embedding of watermarks need to occur simultaneously. When embedding a watermark, we cannot access the complete code generated by the LLM. Thus, we need to make decisions about watermark embedding for the yet-to-be-generated code based on incomplete code, which presents a significant challenge.
3 MCGTest: A prompt dataset for LLM malicious code generation
@@
In this section, we conduct an empirical study using real-world malicious code to help design our watermark. This study results in the creation of MCGTest, a comprehensive dataset of malicious code generation prompts that includes both actual instances of LLM-generated malicious code and potential scenarios.
3.1 Data Collection
@@
To comprehensively collect malicious code generation scenarios with LLMs, we include both real examples and potential use cases.
@@
Existing Instances Collection. This involves gathering existing instances of malicious code generated by LLMs. We collect data from two major technical communities, GitHub [44] and Stack Overflow [45], three main literature databases, Google Scholar [46], arXiv [47], DBLP [48], and the Google search engine [49]. We use the following four keywords for the above six platforms to collect results from January 2023 to March 2024: “large language model malicious code,” “large language model malware,” “GPT malware,” and “GPT malicious code.” In total, we collect $3,644$ results, including code repositories, papers, and articles.
@@
Potential Scenarios Collection. This involves gathering possible scenarios for using LLMs to generate malicious code. For this part, we primarily collect repositories related to malicious code from GitHub. Using the keywords “malicious code” and “malware,” we identify and collect $21,959$ malicious code repositories.
3.2 Data Pre-processing

### v2

# MCGMark: An Encodable and Robust Online Watermark for Tracing LLM-Generated Malicious Code

@@
Abstract.
With the advent of large language models (LLMs), numerous software service providers are developing LLMs tailored for code generation, such as CodeLlama. However, these models can be exploited by malicious developers to generate malicious code, posing severe threats to the software ecosystem. To address this issue, we first conducted an empirical study and built MCGTest, a dataset of $406$ prompts designed to elicit malicious code from LLMs. Leveraging this dataset, we propose MCGMark, a watermarking method to trace and attribute LLM-generated malicious code. MCGMark subtly embeds user-specific information into generated code by controlling the token selection process, ensuring the watermark is imperceptible. Additionally, MCGMark dynamically adjusts the token selection range to induce the LLM to favor high-probability tokens, thus ensuring code quality. Furthermore, by leveraging code structure, MCGMark avoids embedding watermarks into regions easily modified by attackers, such as comments and variable names, enhancing robustness against tampering. Experiments on several advanced LLMs show that MCGMark successfully embeds watermarks in approximately $85\%$ of cases, under the constraint of a $400$-token limit. Moreover, it maintains code quality and demonstrates strong resilience against common code modification. This approach offers a practical solution for tracing malicious code and mitigating the misuse of LLMs.
@@ 1. Introduction
Code generation has become a crucial topic in software engineering (Liu et al., 2023c; Hou et al., 2024). It enables the automatic generation of code snippets from natural language requirements and significantly reduces manual coding efforts (Yu et al., 2024; Jiang et al., 2024). Recently, with the advent and development of large language models (LLMs), their potential in code-related tasks has been widely recognized (Du et al., 2023; Yu et al., 2025). In response, Software Service Providers (SSP) are dedicating efforts to develop LLMs specifically tailored for code generation tasks, such as CodeLlama (Roziere et al., 2023) and DeepSeek-Coder (Guo et al., 2024).
However, despite their benefits, LLMs are also exploited for malicious purposes. Prior research (checkpoint, 2023b; Chen et al., 2024; Madani, 2023) reveals that malicious developers leverage LLMs to develop malware, such as spyware and ransomware. In addition, reports from organizations like CheckPoint (Poireault, 2025) and CrowdStrike (CROWDSTRIKE, 2025) highlight a growing trend of malicious software and cyberattacks facilitated by LLMs. Numerous cases and posts on technical forums further demonstrate the use of LLMs in generating harmful code (tencent, 2023; checkpoint, 2023a; OpenAI, 2023; Land, 2023), posing significant risks to the software ecosystem.
Tracing malicious code generated by LLMs can effectively mitigate the abuse of LLMs. However, the existing methods, such as zero-shot detectors (Mitchell et al., 2023) and fine-tuning language model detectors (Chen et al., 2023), have been proven to be ineffective in practical use (Wu et al., 2023; Yao et al., 2023). For instance, OpenAI discontinued its classifier due to low accuracy (around $26\%$(Land, 2023).
As an alternative, watermarking technology is considered a promising solution for tracing the origin of content generated by LLMs (Fernandez et al., 2023; Takezawa et al., 2023). It embeds identifiable characteristics into the generated content, either explicitly or implicitly, to distinguish and attribute its origin (Li et al., 2023c). However, existing watermarking methods still face several key challenges when applied to tracing malicious code (Li et al., 2023d; Liu et al., 2023b). (1) Traceability and Implicitness. Current watermarking methods primarily focus on detecting whether a piece of content is generated by an LLM, but they overlook tracing the identity of the generator (Lee et al., 2024; Liu et al., 2023a). Moreover, most approaches add watermarks only after the content is generated by the LLM (Chang et al., 2024; Yang et al., 2024a), typically using lexical substitution, which makes them easier to identify and remove (Fernandez et al., 2023). (2) Ensuring code generation quality. Mainstream implicit watermarking techniques are designed for natural language text, whereas code exhibits strong structural constraints and strict syntax requirements, making it challenging to embed watermarks without affecting functionality or quality (He and Vechev, 2023). (3) Resistant to Tampering. Since code elements like comments and variable names can be easily altered by attackers, watermarking methods should account for code structure to improve tamper resistance and robustness (Tipirneni et al., 2024).
@@
To address the aforementioned issues, in this paper, we propose MCGMark, a watermarking framework for tracing LLM-generated code. Our approach implicitly embeds encodable watermarks during the code generation process, considering the code’s structure and ensuring the quality of the generated code. Firstly, MCGMark implicitly embeds encodable watermarks by controlling the LLM’s token selection, ensuring the watermark is difficult to discern while reflecting the generator’s identity. Secondly, MCGMark dynamically obtains the probability distribution of candidate tokens and constrains the LLM’s selection to higher-probability tokens, thereby ensuring the quality of the watermarked code. Lastly, we introduce a watermark skipping mechanism guided by code structure and syntax rules, allowing MCGMark to decide whether to embed watermarks in subsequent code elements during LLM code generation. This ensures that watermarks are not embedded in easily modifiable code elements, such as comments and variable names, thereby enhancing the robustness of the watermark. Additionally, we conduct an empirical study on existing instances of malicious code. We collect $129$ real malicious code examples generated by LLMs and analyze $21,959$ malicious code repositories on GitHub. Based on the empirical study, we construct the first prompt dataset specifically designed for malicious code generation, comprising $406$ tasks to guide watermark design and assess its effectiveness.
MCGMark is designed as a plugin, decoupled from the LLM. During watermark embedding, it requires no additional models, data or tools. During watermark detection, MCGMark does not need to load LLMs. To implement MCGMark, the SSP only needs to adapt its token-matching rules to align with the specific LLM vocabulary. It is important to note that MCGMark, as a watermarking method, cannot eliminate the malicious nature of generated code or prevent its generation. Moreover, MCGMark applies watermarking to all generated code, regardless of its intent. Since benign code can potentially be repurposed to construct malicious software, MCGMark does not attempt to classify code as malicious or benign.
We apply MCGMark to three advanced LLMs to evaluate its effectiveness, while also introducing other baselines for a more comprehensive performance analysis. MCGMark embeds a $24$-bit watermark in $400$ tokens, achieving a watermark embedding success rate of about $85\%$ across different LLMs. Additionally, it outperforms other baselines in watermark detection success rate. Next, we assess the impact of MCGMark on code quality using the CodeBLEU (Ren et al., 2020) and conduct a user study to further validate the results. The results demonstrate that MCGMark achieves significantly higher CodeBLEU scores than baseline methods, confirming its effectiveness in preserving code quality during watermark embedding. Furthermore, we evaluate MCGMark against $500$ program pairs and $1200$ modification attacks, demonstrating its effectiveness in resisting modification attacks. Finally, we analyze the impact of MCGMark’s hyperparameters and evaluate the time overhead of MCGMark.
@@
  • We construct MCGTest, the first dataset for LLM-based malicious code generation, comprising $406$ prompts derived from real-world cases.
@@
  • We propose MCGMark, a robust and encodable watermarking scheme to trace LLM-generated code. MCGMark implicitly embeds user identity information in code generation, ensuring both code quality and robustness against watermark tampering.
  • We evaluate MCGMark on multiple LLMs through comparative experiments. The results show that it successfully embeds a $24$-bit watermark with a success rate of approximately $85\%$ under a $400$ token output limit. Moreover, MCGMark demonstrates competitive performance in preserving code quality, resisting various attacks, and maintaining low time overhead, outperforming existing baseline methods in multiple aspects.
  • We will release the source code of MCGMark and the related datasets after the paper is accepted to support further research.
2. Background and Challenges
2.1. Code Generation of LLM
@@
Large Language Models (LLMs) are large-scale language models based on the Transformer architecture (Bietti et al., 2024) and are trained on massive corpora, typically with billions of parameters or more (Yang et al., 2024b). In recent research, LLMs have demonstrated impressive performance in code generation tasks (Fan et al., 2023) , which can significantly improve software development efficiency (Liu et al., 2024b). The performance of LLMs in code generation tasks has received extensive research attention (Shin et al., 2024; Liu et al., 2024b). Moreover, some Software Service Providers (SSPs) have developed dedicated LLMs specifically designed for code generation, commonly referred to as Code LLMs. Currently, SSPs have developed numerous popular Code LLMs, such as Code Llama (Roziere et al., 2023), DeepSeek-Coder (Guo et al., 2024), and StarCoder2 (Lozhkov et al., 2024).
Figure 1. The motivation example and challenges of design watermark against LLM-generated malicious code.
2.2. Motivation
@@
LLMs are increasingly being exploited by malicious developers for generating malicious code (of Illinois Urbana-Champaign, 2023; Trend, 2023; ThreatDown, 2023). Numerous instances have demonstrated the high efficacy of LLMs in producing harmful software (CROWDSTRIKE, 2025; ThreatDown, 2023; tencent, 2023; checkpoint, 2023a; Trend, 2023). Fig. 1 illustrates a real-world example where an LLM was prompted to generate code for stealing browser history  [^1] https://github.com/AI-Generated-Scripts/GPT-Malware. This irresponsible utilization of LLMs for malicious code generation could pose a significant threat to the security of the software ecosystem.
Moreover, prior studies have demonstrated that LLMs can be easily misused to generate malicious code (Chen et al., 2024; Lin et al., 2024). For instance, RMCBench (Chen et al., 2024) evaluates the resistance of 11 representative LLMs against malicious code generation. The results show that the average refusal rate across all LLMs is only 28.71%, and LLMs with varying generation capabilities can all be used to produce malicious code. Furthermore, malicious developers can employ instruction hijacking (Qiang et al., 2023) and jailbreaking (Niu et al., 2024) to further facilitate the generation of malicious code through LLMs. Therefore, it is imperative to design alternative approaches for LLMs to combat malicious code generation, with watermarking schemes emerging as one of the most promising solutions (Liu et al., 2023b).
2.3. Challenges
@@
Designing watermarks to trace the generation of malicious code introduces several challenges.
  • Traceability & Implicitness. The watermark should be accurately reflect the user’s ID and implicit. Fig.1.(1).a shows an example of a watermark from the work (Lee et al., 2024). Applying this watermark only indicates whether the code was generated by an LLM, failing to trace to a specific user. Fig.1. (1).b illustrates a pattern from the watermarking technique in the work (Li et al., 2024b), which adopts a post-processing watermarking strategy. This technique does not intervene during the code generation process. Instead, it modifies the code after generation, such as performing code transformation (Yang et al., 2024a). However, such approaches rely on predefined transformation patterns, which are inherently limited in applicability and cannot ensure compatibility across diverse code structures. In addition, watermarks based on fixed patterns tend to introduce noticeable artifacts, increasing the risk of being recognized and removed by malicious developers. Therefore, watermarking mechanisms should be designed to remain imperceptible while reliably encoding user-specific information.
@@
  • Ensuring Code Generation Quality. The embedding of watermarks must maintain the quality of the generated code. Fig.1.(2) shows an example of a watermark from the literature (Kirchenbauer et al., 2023a). In this instance, watermark embedding significantly degraded the code quality, rendering the LLM-generated code unusable. In contrast to natural language, code is generally more structured and constrained by strict syntactic and semantic rules (He and Vechev, 2023). Some multi-encoding watermarking techniques attempt to mitigate quality degradation by leveraging the strong generative capabilities of LLMs (Yoo et al., 2023, 2024). However, such approaches are not well-suited for code generation, where even slight modifications to the code may compromise functionality or correctness. Therefore, minimizing the impact of watermarking on code quality during generation remains a key challenge.
@@
  • Resistant to Tampering. Code elements, such as comments and variable names, can be altered without affecting the code’s functionality. If a watermark is embedded in these elements, it can be easily altered or removed. Fig.1.(3).a illustrates a watermark example from the literature (Li et al., 2023d), where the watermark is added to variable names and can be easily modified. Fig.1. (3).b shows an example of a watermark from the literature (Kirchenbauer et al., 2023a), where the watermark is added to comments and can also be easily disrupted. Therefore, the watermark needs to possess sufficient robustness to prevent it from being easily removed. However, online watermark embedding requires that code generation and watermark insertion occur simultaneously. Without access to the complete LLM output during embedding, the watermark must rely on incomplete context, making it difficult to ensure watermark robustness.
3. MCGTest: A prompt dataset for LLM malicious code generation
@@
In this section, we conduct an empirical study on real-world malicious code to help design our watermark. And we conduct MCGTest, a dataset of malicious code generation prompts that includes both actual instances of LLM-generated malicious code and potential scenarios.
3.1. Data Collection
@@
To thoroughly cover malicious code generation scenarios involving LLMs, we include both real-world examples and potential cases to ensure comprehensive coverage.
@@
(Part 1) Existing Instances Collection. This involves gathering existing instances of malicious code generated by LLMs. We collect data from two major technical communities, GitHub (Github, 2024) and Stack Overflow (stackoverflow, 2024), three literature databases, Google Scholar (Google, 2024), arXiv (arXiv, 2024), DBLP (dblp, 2024), and the Google search engine (google, 2024). We use the following four keywords for the above six platforms to collect results from January 2023 to March 2024: “large language model malicious code,” “large language model malware,” “GPT malware,” and “GPT malicious code.” In total, we collect $3,644$ results, including code repositories, papers, and articles.
@@
(Part 2) Potential Scenarios Collection. This involves gathering possible scenarios for using LLMs to generate malicious code. For this part, we primarily collect repositories related to malicious code from GitHub. Using the keywords “malicious code” and “malware,” we identify and collect $21,959$ malicious code repositories.
3.2. Data Pre-processing